OpenStudy (hellokitty17):

A polynomial function can be written as (x + 2)(x + 3)(x - 5). What are the x-intercepts of the graph of this function? (2, 0), (3, 0), (-5, 0) (-2, 0), (-3, 0), (5, 0) (2, 0), (3, 0), (5, 0) (-2, 0), (-3, 0), (-5, 0)

oregonduck (oregonduck):

@Blue_Girl123

oregonduck (oregonduck):

@heretohelpalways

oregonduck (oregonduck):

@ali2x2

OpenStudy (ali2x2):

Split into three equations: x-5 = 0 or x+2 = 0 or x+3 = 0 Look at the first equation: Solve for x. Add 5 to both sides: x = 5 or x+2 = 0 or x+3 = 0 Look at the second equation: Solve for x. Subtract 2 from both sides: x = 5 or x = -2 or x+3 = 0 Look at the third equation: Solve for x. Subtract 3 from both sides: answer would be (-2, 0), (-3, 0), (5, 0)

OpenStudy (hellokitty17):

thanks a lot

OpenStudy (ali2x2):

mhm